OpenStudy (anonymous):

The quadratic function f(x)=px²+qx+r has f(1)=20, and f(2)=11. (Use this information to set up a system) Find the values of the constants p, q, and r. Express f(x) in the form a(x+b)²+c. Use your result to find the smallest value, minimum, of f(x).

OpenStudy (anonymous):

are you sure there is not a piece missing?

OpenStudy (anonymous):

you are being asked to find 3 constants with only two pieces of information you cannot do it or rather, you can find an infinite number of solutions

OpenStudy (anonymous):

The quadratic function f(x)=px²+qx+r has f(0)=35, f(1)=20, and f(2)=11. (Use this information to set up a system) Find the values of the constants p, q, and r. Express f(x) in the form a(x+b)²+c. Use your result to find the smallest value, minimum, of f(x). I forgot the f(0)=35 part.

OpenStudy (anonymous):

lol how did i guess ?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[f(x)=px^2+qx+r\] since \(f(0)=r\) you know \(r=35\)

OpenStudy (anonymous):

now you know \[f(x)=px^2+qx+35\] and we can get the system \[f(1)=p+q+35=20\] or \[p+q=-15\]

OpenStudy (anonymous):

\[f(2)=p\times 2^2+q\times 2+35=11\] so \[4p+2q=-24\]

OpenStudy (anonymous):

solve \[p+q=-15\\ 4p+2q=-24\] to find \(p\) and \(q\)

OpenStudy (anonymous):

Oh, wait, you want me to answer?